OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form.

OpenStudy (anonymous):

answers could be (y + 1) = 3(x – 2) (y – 1) = 3(x + 2) (x + 2) = 3(y – 1) (x – 2) = 3(y + 1)

OpenStudy (anonymous):

@amistre64